Budget   Actual Units produced        10,000    11,050 Materials used in production         &nbs

       Budget

  Actual

Units produced

       10,000

   11,050

Materials used in production

            400 kg

            ?

Material Costs

$      8,000

            ?

Direct Labour (Hours)

      35,000 hrs

    41,010 hrs

Direct Labour Costs

$  385,000

$450,951

Variable Overhead Costs

$  350,000

$411,441

Fixed Overhead Costs

$  160,000

$136,358

Other Information

Overhead is Allocated on Direct Labour Hours

During the year, 602 kg of materials were purchased for $12,000

Beginning Direct Material Inventory:  none

Ending Direct Material Inventory:     39kg

Required:    Calculate the following variances

  1. Material Rate
  2. Material Efficiency
  3. Labour Price
  4. Labour Efficiency
  5. Variable Overhead Rate
  6. Variable Overhead Efficiency
  7. Fixed Overhead Rate
  8. Fixed Overhead Production Volume
Answer #1

A. Material rate variance = Actual cost - Actual Quantity x Standard rate
= $12000 - 602 x $20 = $40 (F)
Standard Rate = $8000 / 400 = $20 per kg

B. Material Efficiency Variance = (Actual Quantity used - Standard Quantity) x Standard Rate
= (563 - 442) x $20 = $2420 (U)
Actual Quantity used = 602-39 = 563
Standard Quantity = 400/10000x11050 = 442

C. Labor Price Variance = Actual Cost - Standard rate x Actual hours
= $450951 - 41010 x $11 = $159 (F)
Standard rate = $385000/35000 = $11 per hour

D. Labor efficiency Variance = (Actual hours - Standard hours) x Standard rate
= (41010 - 38675) x $11 = $25685 (U)
Standard hours = 35000/10000 x 11050 = 38675

As per HOMEWORKLIB RULES we are supposed to solve 4 parts, i have done 4 parts, so kindly post other parts separately